OpenStudy (anonymous):

remove the parentheses ans solve for the variable. 4(4x-3)+4=-3

OpenStudy (anonymous):

and*

OpenStudy (anonymous):

MULTIPLY 4*4X =16X 4*-3=-12 SO 16X-12+4=-3 THEN ADD 12 AND SUBTRACT 4 TO BOTH SIDES IS = TO 16X=11 THE DIVIDE 11/16 AND THATS THE ANSWER

OpenStudy (anonymous):

0.6875=X

OpenStudy (anonymous):

IS IT GOOD NEED HELP IN SOMETHING ELSE OR THERES SOMETHING THAT U DONT UNDERSTAND

OpenStudy (anonymous):

i was getting that and mymathlab was telling me i was wrong :/ hold on i have another

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

4(a-7)+10=2(a-6)

OpenStudy (anonymous):

MULTIPLY 4*A=4A 4*7=28 4A+28+10= 2*A=2A AND 2*-6=-12 SO IS 4A+28+10=2A-12 NOW WE HAVE TO DO LETTER TO THE LEFT AND # TO THE RIGHT SO WE SUBTRACT SO IT IS 2A=50 AND NOW DIVIDE 50/2 = 25 SO A=25

OpenStudy (anonymous):

it's saying that's not right either :(

OpenStudy (anonymous):

ZOORY -50 SO IS A=-25

OpenStudy (anonymous):

ZORRY I DINT PRESS THE MINUS SIGN

OpenStudy (anonymous):

it's saying either way is not right :/ i have worked them out 2 times and got the same as u

OpenStudy (anonymous):

HOW DO HE TEACHES U DONT UNDERSTAND WAIT A SEC

OpenStudy (anonymous):

i send a msj to the good people to check my answer

OpenStudy (anonymous):

Is it saying both of them are incorrect ?

OpenStudy (anonymous):

yes! mymathlab it i got the 1st one wrong cause i ran out of chances for it i have one more for this question and really can't afford to miss it

OpenStudy (anonymous):

Give me the problem and let me see what I can do

OpenStudy (anonymous):

4(a-7)+10=2(a-6)

OpenStudy (mertsj):

16x-12+4=3 Combine like terms: 16x-8=3 Add 8 to both sides 16x=11 Divide both sides by 16 x=11/16

OpenStudy (anonymous):

i got the 1st one right

OpenStudy (anonymous):

yes i did too but for some reason my mathlab is telling me diff.

OpenStudy (anonymous):

4(a - 7) + 10 = 2(a - 6) --- distribute through the parenthesis 4a - 28 + 10 = 2a - 12 --- combine like terms 4a - 18 = 2a - 12 -- subtract 2a from both sides 4a - 2a - 18 = -12 -- add 18 to both sides 4a - 2a = -12 + 18 -- combine like terms 2a = 6 -- divide by 2 a = 3 your 1st problem... 4(4x - 3) + 4 = -3 16x - 12 + 4 = -3 16x - 8 = -3 16x = -3 + 8 16x = 5 x = 5/16

OpenStudy (anonymous):

kelliegirl33 do u mind checking another for me? b4 i post anything on my mathlab?

OpenStudy (anonymous):

sure...what ya got

OpenStudy (anonymous):

solve for the variable 2+y/4=3y/8-4

OpenStudy (anonymous):

2 + (y/4) = 3y/(8-4) --- is this the problem ? or (2 + y)/4 = 3y/(8 - 4)

OpenStudy (anonymous):

the problem had no parentheses in it it's wrote just like i sent u

OpenStudy (anonymous):

is it multiple choice ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

i am getting 48

OpenStudy (anonymous):

(2 + y) / 4 = 3y/(8 - 4) (2 + y) / 4 = 3y/4 this is a proportion, so we will cross multiply (4)(3y) = 4(2 + y) 12y = 8 + y -- subtract y 12y - y = 8 11y = 8 y = 8/11 or 2 + y/4 = 3y/8 - 4 multiply by the LCD, which is 8 16 + 2y = 3y - 32 2y - 3y = -32 - 16 -y = - 48 y = 48 not exactly sure because there is no parenthesis

OpenStudy (anonymous):

i tried 48 with the hopes and we was right thanks 4 ur help!

OpenStudy (anonymous):

sure thing :)